# Prompt-guided Disentangled Representation for Action Recognition

### v1

@@ Abstract
Action recognition is a fundamental task in video understanding. Existing methods typically extract unified features to process all actions in one video, which makes it challenging to model the interactions between different objects in multi-action scenarios. To alleviate this issue, we explore disentangling any specified actions from complex scenes as an effective solution. In this paper, we propose Prompt-guided Disentangled Representation for Action Recognition (ProDA), a novel framework that disentangles any specified actions from a multi-action scene. ProDA leverages Spatio-temporal Scene Graphs (SSGs) and introduces Dynamic Prompt Module (DPM) to guide a Graph Parsing Neural Network (GPNN) in generating action-specific representations. Furthermore, we design a video-adapted GPNN that aggregates information using dynamic weights. Experiments in video action recognition demonstrate the effectiveness of our approach when compared with the state-of-the-art methods. Our code can be found in https://github.com/iamsnaping/ProDA.git.
@@ 1 Introduction
In summary, the major contributions of this work are as follows: 1) We present Prompt-guided Representation Disentanglement for Action Recognition, which disentangles any specified action from multi-action videos, enhancing the interpretability of action understanding; 2) Our method achieves State-of-The-Art performance on multi-label action recognition benchmark. 3) Without requiring action localization supervision, our method demonstrates strong action localization capabilities, further validating its effectiveness and interpretability in understanding complex actions.
@@ 2 Related works
Existing action recognition methods can be broadly categorized into three types. First type is CNN-based methods. To compensate for the lack of temporal modeling in 2D CNNs, two-stream networks introduce an additional temporal stream (e.g., optical flow)[9, 8, 35, 38, 42]. By jointly modeling spatial and temporal features, 3D CNNs are able to better capture spatio-temporal information and thus yield improved performance [36, 4, 43]. The second type is ViT-based methods. Most methods either treat entire frames as input tokens or divide each frame into patches before feeding them into the model [11, 39, 24, 26, 24]. These methods treat the video as a whole, which makes it difficult to model interactions among objects, and often introduces substantial noise. Our work alleviate this by disentangling actions from the SSG, enabling more precise modeling of object interactions. The third type is Objected-centric methods. Object-centric methods construct more compact and efficient video representations by focusing on regions of interest (e.g., objects in the video) [45], and enabling generalization of unseen interaction relationships through structured representations [28].[45, 28, 49, 41, 51]. Nevertheless, these methods remain limited in their ability to effectively model interactions among objects. The incorporation of SSG [51, 17] effectively addresses the difficulty in modeling object interactions. In scene graphs, videos are represented as spatio-temporal graph structures, with objects serving as nodes (e.g., human, cup) and their interactions or spatial-relationship (e.g., holding, above) modeled as edges. [30] processes the scene graph of each frame using a Graph Neural Network (GNN). On the other hand, [40] flattens the scene graphs and feeds them into a ViT [6], leveraging the long-range modeling capability of self-attention [37] to effectively capture interaction transitions. These SSG-based methods still adopt feature extraction approaches that process all actions in a video, which makes it challenging to model interactions between different objects in multi-action scenarios. In contrast, our method disentangles actions from complex scenes, enabling better understanding of the actions
@@ 3.2 Dynamic Prompt Module (DPM)
Action Specification Design. We first represent each action category as a one-hot vector. To specify a combination of actions to disentangle, we construct a multi-hot vector (e.g., $1,0,\dots,1]$), where each $1$ indicates a hot element, i.e., a specified action. By aggregating several one-hot vectors, forming the Action Specification (AS). In designing the AS, we consider two components: one for the actions we intend to disentangle, called the Specified Action Prompt (SAP), and another for the actions we aim to exclude, called the Unspecified Action Prompt (UAP). For the design of SAP, a straightforward approach is to directly input the label-based SAP into the model. However, this can lead to shortcut learning, where the model relies on label cues rather than genuinely parsing the video content. Additionally, in many real-world scenarios, the exact set of actions in a video is not always known. To address these limitations, we further design an alternative version of SAP that includes both the ground-truth actions and distractor actions (i.e., actions not present in the video). This encourages the model to focus on learning discriminative features rather than simply following label hints. For the distractor-injected SAP design, we keep the number of hot elements in the SAP multi-hot vector to $K$, only some of which correspond to ground-truth actions, while the rest are distractors. To prevent the model from exploiting shortcuts, we construct the UAP by inverting the SAP. In this way, the combination of SAP and UAP covers the entire label set, ensuring a balanced and comprehensive specification of both present and absent actions. In certain cases, SAP and UAP may contain only distractors (i.e., no true actions). For such scenarios, we introduce an auxiliary label called "no-action" to explicitly supervise the model.
@@
Action Specification Construction. The AS consists of two types: distractor-injected and non-distractor-injected. The distractor-injected AS includes a combination of present labels (i.e., labels of some actions present in the video) and absent labels (i.e., labels of some actions absent in the video). Enumerating all possible label combinations is computationally infeasible. To address this, we adopt a random sampling strategy. For a video with $L$ ground-truth labels, the distractor-injected SAPs, we generate $L+1$ different multi-hot vectors with a length $K$. Among them, $i$-th SAP contains $i$ random sampled present labels, and KiK-i random sampled absent labels. In contrast, the non-distractor-injected SAPs, we generate $L$ different multi-hot vectors, only contains random sampled present labels. UAPs are then obtained by inverting each SAP. This results in $2L+1$ different pairs of SAP and UAP. This design enables the model to learn from a diverse set of action combinations, promoting robust disentanglement by exposing it to both clean and noisy prompts while maintaining efficiency. For better clarity, we include illustrative examples in the supplementary material.
@@ 3.3 Video Graph Parsing Neural Network (VGPNN)
Video Graph Parsing Neural Network (VGPNN). VGPNN is designed to disentangle features by decomposing a SSG into two subgraphs, guided by a given AS. To achieve selective information aggregation, with the same to GPNN [31], each VGPNN layer processes the entire SSG in three sequential steps. In Link step, we perform temporal modeling of edges $e^{j,k}=\{e_{1}^{j,k},e_{i}^{j,k},\dots,e_{N}^{j,k}\}$ via self-attention [37] over the same interaction across time. In Message step, we generate dynamic weights from the edge features, which are then multiplied with the messages, where each message is composed of a node’s features and its edge features. In Update step, we update all nodes by adding the weighted messages to the corresponding nodes directly. A MLP is then used to fuse the updated and original node features.
@@ 4.1 Experiment Setup
Dataset. The Charades dataset [34] consists of 9848 videos, each with an average duration of 30 seconds. It covers 157 action categories, with each video containing an average of 6.8 distinct actions. Multiple actions can occur simultaneously, making the recognition task particularly challenging. Based on Charades, the Action Genome dataset [17] provides fine-grained annotations by decomposing actions and focusing on video clips where the actions take place. It contains 234K keyframes, with annotations for 476K object bounding boxes and 1.72 million object relationships. In the multi-label action recognition setting, the input is a video sequence, and the goal is to predict multiple action labels associated with it. Following the protocol of [40, 17], we evaluate multi-label action recognition performance on the Charades dataset using the Mean Average Precision (mAP).
Experiment Setup. All experiments are conducted under the same design. For each video, the network takes $N$ frames as input. During training, frames are randomly sampled from the video, while during validation, the frames are extracted evenly over the whole video. In the first stage, the entire network is trained end-to-end, and the resulting weights are used to initialize the second stage. In the second stage, only the classifiers for $a_{u},a_{s},a_{t}$ are trained. Additionally, to enable deeper analysis, we introduce an auxiliary prediction $a_{m}$, which is inferred from the features before disentanglement (features from Video Encoder as illustrated in Figure 2). Its corresponding classifier is also trained during the second stage.
@@ 4.2 Ablation Study
model pretrain Modality mAP model pretrain Modality mAP
SGFB+NL [17] K400 RGB 44.3 SGFB+NL [17] K400 RGB+SSG 60.3
I3D+NL [4] K400 RGB 37.5 LaIAR [40] K400 RGB+Bbox 63.6
OR2G+NL [30] K400 RGB 44.9 OR2G [30] K400 RGB+SSG 63.3
LaIAR+NL [40] K400 RGB 45.1 LaIAR+NL [40] K400 RGB+Bbox 67.4
VicTR(B/16) [20] CLIP RGB 50.1 OR2G+NL [30] K400 RGB+SSG 67.5
VicTR(L/14) [20] CLIP RGB 57.6 Ours(B/16) CLIP RGB+SSG 71.1
4.3 Compared with State-of-the Art Methods
We compare the action recognition accuracy of the proposed method and the state-of-the-art methods (SoTA) on the Charades. Table 4 presents our results on the Charades dataset. ProDA achieves state-of-the-art performance. Our method outperforms several CNN- and ViT-based approaches, including I3D [4] and VicTR [20]. This demonstrates that our approach can better capture object interactions within structured data (e.g., SSG), enabling a deeper understanding of video. Notably, our method with a CLIP-B/16 [32] still surpasses VicTR (L/14) [20], even though we only use an image encoder, while VicTR [20] additionally employs a text encoder. Furthermore, our method also outperforms several approaches that leverage SSG [30, 40]. This demonstrates that explicitly disentangling and processing actions separately leads to a more accurate understanding of complex actions. It is also worth noting that our method achieves this performance without requiring any additional external models.
4.4 Further Analysis
Disentanglement Robustness Analysis. To evaluate the robustness of our action disentanglement, we analyze the $a_{s}$ accuracy of videos with different numbers of ground-truth labels (denoted as label nums in Table 6 6) when exposed to varying numbers of present labels in the SAP, under both distractor-injected and non-distractor settings. As shown in Table 6, in the distractor-injected case (SAP contains the labels not present in video), we observe a non-monotonic trend: accuracy drops when the number of present labels increases from 5 to 12, but rises again as present labels dominate the input. We attribute the initial drop to the increasing complexity of the prediction task due to more competing present labels under noisy input. The subsequent performance increase is attributed to reduced noise as more present labels are included. Despite the precision varying between 85% and 100% under distractor-injected, our model demonstrates robust performance across different input conditions. Even in the worst-case scenario, where the precision is at its lowest (85%), the model still maintains a relatively high accuracy. This indicates its ability to handle incomplete or noisy data effectively. Moreover, in the non-distractor-injected case (SAPs only contain labels present in video), as shown in Table 6, the model achieves nearly perfect accuracy, reaching close to 100%. This shows that in ideal conditions, the model is able to fully exploit the available information. The robustness in noisy conditions and strong performance on clean data indicate that our model generalizes well and benefits from mutual reinforcement between the two settings.
@@ 4.4 Further Analysis
Temporal Action Localization. Although our method does not explicitly design a localization module, the disentangled representation naturally provides cues for action localization. Specifically, within our readout function, we aggregate temporal features using learned frame-level weights $s$. These weights reflect the relevance of each frame to the target action, and can thus be interpreted as frame-level localization scores. By normalizing these weights, we obtain a per-frame prediction that serves as our localization output. To obtain frame-level ground truth for localization, we project the action segment annotations onto frames: all frames within the annotated action duration are labeled as positive (1), and the rest as negative (0). We binarize the normalized weights $s$ using a threshold $\theta$, assigning 1 to frames with $s>\theta$ and 0 otherwise, and evaluate the resulting predictions using frame-level mAP based on $a_{s}$. As shown in Table 8 8, we observe that the variation in average precision across different $\theta$ values is larger than the variation across different IoU thresholds, and the gap between distractor-injected and non-distractor-injected settings narrows as $\theta$ increases. This indicates that the introduced noise (i.e., absent labels) mainly adds irrelevant information without significantly disrupting the separation of true actions, demonstrating the robustness of our method. In the non-distractor-injected setting, the model achieves notably better localization performance, confirming that the high recognition accuracy is not due to shortcut learning, but rather due to effective utilization of SAP information for representation disentanglement under clean inputs.

### v3

@@ Abstract
Action recognition is a fundamental task in video understanding. Existing methods typically extract unified features to process all actions in one video, which makes it challenging to model the interactions between different objects in multi-action scenarios. To alleviate this issue, we explore disentangling any specified actions from complex scenes as an effective solution. In this paper, we propose Prompt-guided Disentangled Representation for Action Recognition (ProDA), a novel framework that disentangles any specified actions from a multi-action scene. ProDA leverages Spatio-temporal Scene Graphs (SSGs) and introduces Dynamic Prompt Module (DPM) to guide a Graph Parsing Neural Network (GPNN) in generating action-specific representations. Furthermore, we design a video-adapted GPNN that aggregates information using dynamic weights. Extensive experiments on two complex video action datasets, Charades and SportsHHI, demonstrate the effectiveness of our approach against state-of-the-art methods. Our code can be found in https://github.com/iamsnaping/ProDA.git.
@@ 1 Introduction
In summary, the major contributions of this work are as follows: 1) We present a novel framework ProDA, which disentangles any specified action from multi-action videos, enhancing the interpretability of action understanding; 2) Our method achieves State-of-The-Art performance on multi-label action recognition benchmark. 3) Without requiring action localization supervision, our method demonstrates strong action localization capabilities, further validating its effectiveness and interpretability in understanding complex actions.
@@ 2 Related works
Existing action recognition methods can be broadly categorized into three types. First type is CNN-based methods. To compensate for the lack of temporal modeling in 2D CNNs, two-stream networks introduce an additional temporal stream (e.g., optical flow)[10, 9, 41, 44, 49]. By jointly modeling spatial and temporal features, 3D CNNs are able to better capture spatio-temporal information and thus yield improved performance [42, 4, 51]. The second type is ViT-based methods. Most methods either treat entire frames as input tokens or divide each frame into patches before feeding them into the model [13, 45, 27, 30, 27]. These methods treat the video as a whole, which makes it difficult to model interactions among objects, and often introduces substantial noise. Our work alleviate this by disentangling actions from the SSG, enabling more precise modeling of object interactions. The third type is Objected-centric methods. Object-centric methods construct more compact and efficient video representations by focusing on regions of interest (e.g., objects in the video) [53], and enabling generalization of unseen interaction relationships through structured representations [53, 33, 57, 48, 59]. Nevertheless, these methods remain limited in their ability to effectively model interactions among objects. The incorporation of SSG [59, 20] effectively addresses the difficulty in modeling object interactions. In scene graphs, videos are represented as spatio-temporal graph structures, with objects serving as nodes (e.g., human, cup) and their interactions or spatial-relationship (e.g., holding, above) modeled as edges. [35] processes the scene graph of each frame using a Graph Neural Network (GNN). On the other hand, [47] flattens the scene graphs and feeds them into a ViT [7], leveraging the long-range modeling capability of self-attention [43] to effectively capture interaction transitions. These SSG-based methods still adopt feature extraction approaches that process all actions in a video, which makes it challenging to model interactions between different objects in multi-action scenarios. In contrast, our method disentangles actions from complex scenes, enabling better understanding of the actions
@@ 3.2 Dynamic Prompt Module (DPM)
Action Specification Design. We first represent each action category as a one-hot vector. To specify a combination of actions to disentangle, we construct a multi-hot vector (e.g., $1,0,\dots,1]$), where each $1$ indicates a hot element, i.e., a specified action. By aggregating several one-hot vectors, forming the Action Specification (AS). In designing the AS, we consider two components: one for the actions we intend to disentangle, called the Specified Action Prompt (SAP), and another for the actions we aim to exclude, called the Unspecified Action Prompt (UAP). For the design of SAP, a straightforward approach is to directly input the label-based SAP into the model. However, this can lead to shortcut learning, where the model relies on label cues rather than genuinely parsing the video content. Additionally, in many real-world scenarios, the exact set of actions in a video is not always known. To address these limitations, we further design an alternative version of SAP that includes both the ground-truth actions and distractor actions (i.e., actions not present in the video). This encourages the model to focus on learning discriminative features rather than simply following label hints. For the distractor-injected SAP design, we keep the number of hot elements in the SAP multi-hot vector to $K$, only some of which correspond to ground-truth actions, while the rest are distractors. To prevent the model from exploiting shortcuts, we construct the UAP by inverting the SAP. In this way, the combination of SAP and UAP covers the entire label set, ensuring a balanced and comprehensive specification of both present and absent actions (For better clarity, we include illustrative examples in Sec. C.1). In certain cases, SAP and UAP may contain only distractors (i.e., no true actions). For such scenarios, we introduce an auxiliary label called "no-action" to explicitly supervise the model.
@@
Action Specification Construction. The AS consists of two types: distractor-injected and non-distractor-injected. The distractor-injected AS includes a combination of present labels (i.e., labels of some actions present in the video) and absent labels (i.e., labels of some actions absent in the video). Enumerating all possible label combinations is computationally infeasible. To address this, we adopt a random sampling strategy. For a video with $L$ ground-truth labels, the distractor-injected SAPs, we generate $L+1$ different multi-hot vectors with a length $K$. Among them, $i$-th SAP contains $i$ random sampled present labels, and KiK-i random sampled absent labels. In contrast, the non-distractor-injected SAPs, we generate $L$ different multi-hot vectors, only contains random sampled present labels. UAPs are then obtained by inverting each SAP. This results in $2L+1$ different pairs of SAP and UAP. This design enables the model to learn from a diverse set of action combinations, promoting robust disentanglement by exposing it to both clean and noisy prompts while maintaining efficiency. For better clarity, we include illustrative examples in Sec. C.2.
@@ 3.3 Video Graph Parsing Neural Network (VGPNN)
Video Graph Parsing Neural Network (VGPNN). VGPNN is designed to disentangle features by decomposing a SSG into two subgraphs, guided by a given AS. To achieve selective information aggregation, with the same to GPNN [37], each VGPNN layer processes the entire SSG in three sequential steps. In Link step, we perform temporal modeling of edges $e^{j,k}=\{e_{1}^{j,k},e_{i}^{j,k},\dots,e_{N}^{j,k}\}$ via self-attention [43] over the same interaction across time. In Message step, we generate dynamic weights from the edge features, which are then multiplied with the messages, where each message is composed of a node’s features and its edge features. In Update step, we update all nodes by adding the weighted messages to the corresponding nodes directly. A MLP is then used to fuse the updated and original node features. More details please refer to Sec. B.
@@ 3.4 Optimization
model pretrain Modality mAP model pretrain Modality mAP
SGFB+NL [20] K400 RGB 44.3 SGFB+NL [20] K400 RGB+SSG 60.3
I3D+NL [4] K400 RGB 37.5 LaIAR [47] K400 RGB+Bbox 63.6
OR2G+NL [35] K400 RGB 44.9 OR2G [35] K400 RGB+SSG 63.3
LaIAR+NL [47] K400 RGB 45.1 LaIAR+NL [47] K400 RGB+Bbox 67.4
VicTR(B/16) [23] CLIP RGB 50.1 OR2G+NL [35] K400 RGB+SSG 67.5
VicTR(L/14) [23] CLIP RGB 57.6 Ours(B/16) CLIP RGB+SSG 71.1
Table 2: Human-Human Interaction Classification results in SportsHHI. SportsHHI is our reimplementation based on the official SportsHHI code, with minor modifications (e.g., parameter size) to match the capacity of ProDA for fair comparison.
method mAP R@150 R@100 R@50 R@20 Params
STTran [5] 3.31 71.29 62.91 42.67 22.14 -
HORT  [19] 3.75 78.57 67.78 50.33 26.96 -
SlowFast [11] 5.00 81.66 74.00 52.74 26.82 -
ACARN  [36] 5.44 82.85 75.22 56.53 31.77 -
SportHHI  [50] (baseline) 10.69 89.25 82.93 68.13 43.72 111.18M
SportsHHI (reimpletmented) 10.16 ($\downarrow$0.53) 90.07 ($\uparrow$0.82) 84.83 ($\uparrow$1.9) 68.40 ($\uparrow$0.37) 40.16 ($\downarrow$ 3.56) 123.11M
Ours (baseline+ProDA) 13.55 ($\uparrow$2.86) 92.54 ($\uparrow$3.29) 87.91 ($\uparrow$4.98) 72.69 ($\uparrow$4.56) 47.81 ($\uparrow$4.09) 123.19M
@@ 4.1 Experiment Setup
Dataset. (1) The Charades dataset [40] consists of 9,848 videos, each with an average duration of 30 seconds. It covers 157 action categories, with each video containing an average of 6.8 distinct actions. Multiple actions can occur simultaneously, making the recognition task particularly challenging. Based on Charades, the Action Genome dataset [20] provides fine-grained annotations by decomposing actions and focusing on video clips where the actions take place. It contains 234K keyframes, with annotations for 476K object bounding boxes and 1.72 million object relationships. (2) The SportsHHI dataset [50] is a specially designed dataset focusing on human-human interaction in sports. It covers basketball and volleyball games, and video data are selected from MultiSports [28]. This dataset labels 34 interaction classes with 118,075 human bounding boxes and 50,649 interaction instances. Following [47, 50], we conduct evaluations on Charades for multi-label action recognition and on SportsHHI for Human-Human interaction Classification (HHICls), adopting mean Average Precision (mAP) as the metric.
Experiment Setup. (1) On Charades, all experiments are conducted under the same design. For each video, the network takes $N$ frames as input. During training, frames are randomly sampled from the video, while during validation, the frames are extracted evenly over the whole video. In the first stage, the entire network is trained end-to-end, and the resulting weights are used to initialize the second stage. In the second stage, only the classifiers for $a_{u},a_{s},a_{t}$ are trained. Additionally, to enable deeper analysis, we introduce an auxiliary prediction $a_{m}$, which is inferred from the features before disentanglement (features from Video Encoder as illustrated in Figure 2). Its corresponding classifier is also trained during the second stage. (2) On SportsHHI, we adopt the official open-source model and training code without altering any training settings. For fair evaluation, we conduct two experiments: 1) extending the official baseline with our proposed modules, namely the DPM and AD Loss; and 2) constructing a parameter-matched baseline by enlarging its model size to match ours, without incorporating any proposed components.
4.2 Compared with State-of-The Art Methods
Charades. We compare the action recognition accuracy of the proposed method and the State-of-The-Art methods (SoTA) on the Charades. Table 1 presents our results on the Charades dataset. ProDA achieves state-of-the-art performance. Our method outperforms several CNN- and ViT-based approaches, including I3D [4] and VicTR [23]. This demonstrates that our approach can better capture object interactions within structured data (e.g., SSG), enabling a deeper understanding of video. Notably, our method with a CLIP-B/16 [38] still surpasses VicTR (L/14) [23], even though we only use an image encoder, while VicTR [23] additionally employs a text encoder. Furthermore, our method also outperforms several approaches that leverage SSG [35, 47]. This demonstrates that explicitly disentangling and processing actions separately leads to a more accurate understanding of complex actions. It is also worth noting that our method achieves this performance without requiring any additional external models.
SportsHHI. We compare our method with the SoTA baseline and its parameter-increased variant, as reported in Table 2. Enlarging the baseline leads to a 10.8% increase in parameters, whereas our proposed DPM introduces only 1.46M additional parameters (1.18% of the total). Despite such a small overhead, our method achieves notable improvements across all metrics. Even under comparable parameter settings, it consistently outperforms the baseline. These results demonstrate the strong generalization capability of our approach and confirm that the performance gain primarily arises from the proposed AD Loss rather than from increased model capacity.
@@ 4.3 Ablation Study
4.4 Further Analysis
Disentanglement Robustness Analysis. To evaluate the robustness of our action disentanglement, we analyze the $a_{s}$ accuracy of videos with different numbers of ground-truth labels (denoted as label nums in Table 7 7) when exposed to varying numbers of present labels in the SAP, under both distractor-injected and non-distractor settings. As shown in Table 7, in the distractor-injected case (SAP contains the labels not present in video), we observe a non-monotonic trend: accuracy drops when the number of present labels increases from 5 to 12, but rises again as present labels dominate the input. We attribute the initial drop to the increasing complexity of the prediction task due to more competing present labels under noisy input. The subsequent performance increase is attributed to reduced noise as more present labels are included. Despite the precision varying between 85% and 100% under distractor-injected, our model demonstrates robust performance across different input conditions. Even in the worst-case scenario, where the precision is at its lowest (85%), the model still maintains a relatively high accuracy. This indicates its ability to handle incomplete or noisy data effectively. Moreover, in the non-distractor-injected case (SAPs only contain labels present in video), as shown in Table 7, the model achieves nearly perfect accuracy, reaching close to 100%. This shows that in ideal conditions, the model is able to fully exploit the available information. The robustness in noisy conditions and strong performance on clean data indicate that our model generalizes well and benefits from mutual reinforcement between the two settings.
Temporal Action Localization. Although our method does not explicitly design a localization module, the disentangled representation naturally provides cues for action localization. Specifically, within our readout function, we aggregate temporal features using learned frame-level weights $s$. These weights reflect the relevance of each frame to the target action, and can thus be interpreted as frame-level localization scores. By normalizing these weights, we obtain a per-frame prediction that serves as our localization output. To obtain frame-level ground truth for localization, we project the action segment annotations onto frames: all frames within the annotated action duration are labeled as positive (1), and the rest as negative (0). We binarize the normalized weights $s$ using a threshold $\theta$, assigning 1 to frames with $s>\theta$ and 0 otherwise, and evaluate the resulting predictions using frame-level mAP based on $a_{s}$. As shown in Table 9 9, we observe that the variation in average precision across different $\theta$ values is larger than the variation across different IoU thresholds, and the gap between distractor-injected and non-distractor-injected settings narrows as $\theta$ increases. This indicates that the introduced noise (i.e., absent labels) mainly adds irrelevant information without significantly disrupting the separation of true actions, demonstrating the robustness of our method. In the non-distractor-injected setting, the model achieves notably better localization performance, confirming that the high recognition accuracy is not due to shortcut learning, but rather due to effective utilization of SAP information for representation disentanglement under clean inputs.
Figure 4: t-SNE visualization of disentangled action features across 157 classes on Charades [40]. (a) Training set. (b) Test set.